# Inferring resource competition in microbial communities from time series

### v1

@@ II.7 Applying temporal metrics to marine data
We use the time series of the relative abundances of the $N=31$ eukaryotic OTUs to compute the four pairwise metrics as described in the previous sections: the equal-time correlation $\mathbf{C}(0)$ its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$. Ideally, we would like to compare the pairwise metrics with the elements of the resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$. However, since there is no a priori measurement or knowledge of the guild structure in natural communities, we use phylogenetic distance as a proxy, following previous work sireci_environmental_2023 . The intuition is that the closer the species are genetically, the more likely they are to share the resources martiny_phylogenetic_2013 .
@@ IV.2 Finite size scaling in ECRMs
In order to bridge the toy model with a small number of species and guilds with realistic ecological systems with large numbers of species and guilds, we scaled the parameters of the ECRM given different system sizes. As we increase the number of species $N$, we want to also increase both the number of guilds $k$ and the number of species within each guild $N/k$. Hence, we set the number of guilds $k\propto\sqrt{N}$. The number of resources should also increase as $M\propto N$. Setting the probability of species $i$ to have a non-zero intake of its preferred resource as $p_{\text{in}}$, and the probability of species $i$ to have a non-zero intake of its non-preferred resource as $p_{\text{out}}$, we can compute the expectation value of the average degree of each species in the effective species-species interaction due to resource utilization overlap, the expectation value of $(CC^{\intercal})_{ij}$, etc.
@@
$P(CC^{\intercal}_{ij}=0)=\left(1-p_{\text{in}}^{2}\right)^{M/k}\left(1-p_{\text{out}}^{2}\right)^{M(1-1/k)}.$ (16)
@@
$P(CC^{\intercal}_{ij}=0)=\left(1-p_{\text{in}}p_{\text{out}}\right)^{2M/k}\left(1-p_{\text{out}}^{2}\right)^{M(1-2/k)}.$ (17)
@@
For proper scaling, we want that, in the limit of large $N$, the average degree $d$ converges to a constant value. The choice $p_{\text{in}}\propto N^{-1/2}$ and $p_{\text{out}}\propto N^{-1}$ satisfies this condition, by ensuring that the average degrees towards in-guild and out-of-guild species are both finite. Notice that this scaling results in a $1/N$ scaling for both the expectation value and the variance of elements of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix.
@@ Acknowledgments
Figure S2: Coherence and CPSD outperforms equal-time correlation measures in recovering the resource-utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$ for all sizes of systems. (a) Spearman’s rank correlation coefficient between the four considered pairwise observables (equal time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$) and the elements of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix, for three families of ECRM systems with $N=6,24,96$ defined in Table 1. Error bars represent standard deviation across 100 random realizations of species abundances for systems with $N=6$ and with $N=24$ (10 random ECRM systems each with 10 random OU environmental drives with $\omega_{\alpha}=1$), and across 10 random realizations of species abundances for systems with $N=96$ (10 random ECRM systems each with 1 random OU environmental drive with $\omega_{\alpha}=1$). (b) Area under curve (AUC) for link detection for the same systems as in (a). Each AUC is computed for the same ECRM system, averaged over the ROC of all 10 random OU drives. (c) AUC for guild detection after single linkage clustering for the same systems as in (a).
@@
Figure S3: Influence of finite time-series duration and sampling rate on Spearman’s correlation coefficient between the pairwise measures of species abundance and the resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$. Spearman’s correlation coefficient (colorbar) between elements of the $\mathbf{C}\mathbf{C}^{\intercal}$ matrix and elements of (a) Equal-time correlation $\mathbf{C}(0)$, (b) its absolute value $\lvert\mathbf{C}(0)\rvert$, (c) the total absolute value of CPSD, and (d) the total coherence, as the function of sampling duration $t_{f}$ and the sampling interval $\Delta t_{\rm sampling}$. The ECRM is the same as the one considered in Fig. 3 with $N=6$, with OU environmental drive with $\omega_{\alpha}=1$ and guild-structure bias $q=0.4$.
@@
Figure S4: Link detection and guild detection of the resource utilization overlap matrix $\mathbf{C}\mathbf{C}^{\intercal}$ using pairwise measurements of the relative abundance (instead of absolute abundance) of species concentration, for simulated ECRM ecosystems of different sizes (Table 1), with guild-structure bias $q=0.5$. (a) Effective resource-utilization matrix $\mathbf{C}\mathbf{C}^{\intercal}$ with $N=6$ species, $M=30$ resources, $k=2$ guilds, and $q=0.5$. (b) ROC curve of link detection for the four pairwise observables. Legend for the colors is given next to panel (c). (d) Same as (c) but for guild detection, after performing single linkage clustering on the pairwise metrics with a tunable threshold. (e-h) Same as (a-d) but for $N=24$, $M=120$, and $k=4$. (i-l) Same as (a-d) but for $N=96$, $M=480$, and $k=8$. For all ECRMs, the environment is driven by OU processes with the intrinsic timescale set to $\omega_{\alpha}=1$. For all system sizes, the total duration of the simulated trajectory is set to $t_{f}=20000$, and the sampling timestep is set to $\Delta t_{\text{sampling}}=1$.

### v2

@@ II.7 Applying temporal metrics to marine data
We use the time series of the relative abundances of the $N=31$ eukaryotic OTUs to compute the four pairwise metrics as described in the previous sections: the equal-time correlation $\mathbf{C}(0)$ its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$. Ideally, we would like to compare the pairwise metrics with the elements of the resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$. However, since there is no a priori measurement or knowledge of the guild structure in natural communities, we use phylogenetic distance as a proxy, following previous work sireci_environmental_2023 . The intuition is that the closer the species are genetically, the more likely they are to share the resources martiny_phylogenetic_2013 .
@@ IV.2 Finite size scaling in ECRMs
In order to bridge the toy model with a small number of species and guilds with realistic ecological systems with large numbers of species and guilds, we scaled the parameters of the ECRM given different system sizes. As we increase the number of species $N$, we want to also increase both the number of guilds $k$ and the number of species within each guild $N/k$. Hence, we set the number of guilds $k\propto\sqrt{N}$. The number of resources should also increase as $M\propto N$. Setting the probability of species $i$ to have a non-zero intake of its preferred resource as $p_{\text{in}}$, and the probability of species $i$ to have a non-zero intake of its non-preferred resource as $p_{\text{out}}$, we can compute the expectation value of the average degree of each species in the effective species-species interaction due to resource utilization overlap, the expectation value of $(\mathbf{G}\mathbf{G}^{\intercal})_{ij}$, etc.
@@
$P((\mathbf{G}\mathbf{G}^{\intercal})_{ij}=0)=\left(1-p_{\text{in}}^{2}\right)^{M/k}\left(1-p_{\text{out}}^{2}\right)^{M(1-1/k)}.$ (16)
@@
$P((\mathbf{G}\mathbf{G}^{\intercal})_{ij}=0)=\left(1-p_{\text{in}}p_{\text{out}}\right)^{2M/k}\left(1-p_{\text{out}}^{2}\right)^{M(1-2/k)}.$ (17)
@@
For proper scaling, we want that, in the limit of large $N$, the average degree $d$ converges to a constant value. The choice $p_{\text{in}}\propto N^{-1/2}$ and $p_{\text{out}}\propto N^{-1}$ satisfies this condition, by ensuring that the average degrees towards in-guild and out-of-guild species are both finite. Notice that this scaling results in a $1/N$ scaling for both the expectation value and the variance of elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix.
@@ IV.4 Statistical tests for binned pairwise metrics against index shuffled null model
Code availability
Codes and data generating all figures in this manuscript are available on OSF (https://doi.org/10.17605/osf.io/hvnjt).
@@ Acknowledgments
Figure S2: Coherence and CPSD outperforms equal-time correlation measures in recovering the resource-utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ for all sizes of systems. (a) Spearman’s rank correlation coefficient between the four considered pairwise observables (equal time correlation $\mathbf{C}(0)$ and its absolute value $\lvert\mathbf{C}(0)\rvert$, the total magnitude of CPSD $\mathcal{P}$, and the total coherence $\mathcal{TC}$) and the elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix, for three families of ECRM systems with $N=6,24,96$ defined in Table 1. Error bars represent standard deviation across 100 random realizations of species abundances for systems with $N=6$ and with $N=24$ (10 random ECRM systems each with 10 random OU environmental drives with $\omega_{\alpha}=1$), and across 10 random realizations of species abundances for systems with $N=96$ (10 random ECRM systems each with 1 random OU environmental drive with $\omega_{\alpha}=1$). (b) Area under curve (AUC) for link detection for the same systems as in (a). Each AUC is computed for the same ECRM system, averaged over the ROC of all 10 random OU drives. (c) AUC for guild detection after single linkage clustering for the same systems as in (a).
@@
Figure S3: Influence of finite time-series duration and sampling rate on Spearman’s correlation coefficient between the pairwise measures of species abundance and the resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$. Spearman’s correlation coefficient (colorbar) between elements of the $\mathbf{G}\mathbf{G}^{\intercal}$ matrix and elements of (a) Equal-time correlation $\mathbf{C}(0)$, (b) its absolute value $\lvert\mathbf{C}(0)\rvert$, (c) the total absolute value of CPSD, and (d) the total coherence, as the function of sampling duration $t_{f}$ and the sampling interval $\Delta t_{\rm sampling}$. The results are averaged over 50 random realizations of the ECRMs with $N=6$, with OU environmental drive with $\omega_{\alpha}=1$ and guild-structure bias $q=0.4$.
@@
Figure S4: Link detection and guild detection of the resource utilization overlap matrix $\mathbf{G}\mathbf{G}^{\intercal}$ using pairwise measurements of the relative abundance (instead of absolute abundance) of species concentration, for simulated ECRM ecosystems of different sizes (Table 1), with guild-structure bias $q=0.5$. (a) Effective resource-utilization matrix $\mathbf{G}\mathbf{G}^{\intercal}$ with $N=6$ species, $M=30$ resources, $k=2$ guilds, and $q=0.5$. (b) ROC curve of link detection for the four pairwise observables. Legend for the colors is given next to panel (c). (d) Same as (c) but for guild detection, after performing single linkage clustering on the pairwise metrics with a tunable threshold. (e-h) Same as (a-d) but for $N=24$, $M=120$, and $k=4$. (i-l) Same as (a-d) but for $N=96$, $M=480$, and $k=8$. For all ECRMs, the environment is driven by OU processes with the intrinsic timescale set to $\omega_{\alpha}=1$. For all system sizes, the total duration of the simulated trajectory is set to $t_{f}=20000$, and the sampling timestep is set to $\Delta t_{\text{sampling}}=1$.